OpenStudy (anonymous):

Help please?! Been stuck on this one question for hours. A jelly donut that weighs 1.5 tons and has a circumference of 50 ft what's the diameter? (C=ND) Thank you!

OpenStudy (lgbasallote):

what is that C = ND? what do the letters mean?

OpenStudy (anonymous):

Don't know, one of the reasons I'm stuck. :P

OpenStudy (lgbasallote):

if you only need diameter then the mass is not important...just use circumference = pi x diameter so d = C/pi d = 50/pi d = 15.92 ft...is that in the choices? btw...that's one big donut

OpenStudy (nottim):

I could get a meal.

OpenStudy (anonymous):

It's either 12ft, 10ft, 14ft, or 16ft.

OpenStudy (nottim):

I think the weight is just a red herring (that just opinion).

OpenStudy (lgbasallote):

16 ft if you round it off

OpenStudy (anonymous):

Positive?

OpenStudy (nottim):

I'm 80%.

OpenStudy (lgbasallote):

yes. my solution proves my point! :D

OpenStudy (nottim):

Now I'm batman%.

OpenStudy (lgbasallote):

that's a lot...

OpenStudy (nottim):

Yes. Above human-recognizable perfection? Obviously.

OpenStudy (lgbasallote):

im sherlock % so i beat you ha!

OpenStudy (nottim):

Let's discuss this in chat. I have so many counterarguments.

OpenStudy (nottim):

Icelandic bagel probably wouldn't care for them all.